OpenStudy (barrelracing):

I will fan and medal. Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 3:35 to 3:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? You must show all of your work.

OpenStudy (anonymous):

how many minutes in an hour?

OpenStudy (barrelracing):

60

OpenStudy (anonymous):

and how many degrees in a circle?

OpenStudy (barrelracing):

360

OpenStudy (anonymous):

so how may degrees per minute?

OpenStudy (barrelracing):

6

OpenStudy (anonymous):

actually since your question is given in radians, we can forget about degrees

OpenStudy (anonymous):

or, now that you have 6 degrees, we can convert the 6 degrees to radians do you know how to do that?

OpenStudy (barrelracing):

i dont know how to do that

OpenStudy (anonymous):

put it over \(180\) and multply by \(\pi\) i.e.. \[6^\circ=\frac{6\pi}{180}\] which you can reduce

OpenStudy (anonymous):

i mean reduce the fraction, leave the \(\pi\) there

OpenStudy (barrelracing):

so for number 1 its 30?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

on minute is \[\frac{\pi}{30}\] radians

OpenStudy (anonymous):

for 20 minutes, multiply that by 20

OpenStudy (barrelracing):

im confused what i multiply by 20

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

is it clear that from 3:35 to 3:55 is 20 minutes?

OpenStudy (barrelracing):

yes

OpenStudy (anonymous):

and it is also clear that 20 minutes is one third of an hour?

OpenStudy (barrelracing):

yes

OpenStudy (anonymous):

and do you know that the entire circle, one complete rotation, is an angle of \(2\pi\)?

OpenStudy (barrelracing):

yes

OpenStudy (anonymous):

so your 20 minutes is one third of that

OpenStudy (barrelracing):

so what should i type for the first one? I put there is 20 minutes in between

OpenStudy (anonymous):

i have no idea what you should type in for what the answer to the first question "How many radians does the minute hand move from 3:35 to 3:55? " is \[\large \frac{2\pi}{3}\]

OpenStudy (barrelracing):

ok

OpenStudy (anonymous):

since twenty minutes is one third of 60 minutes, and two pi divided by 3 is what i wrote above

OpenStudy (barrelracing):

thank you. i will be right back let me type this up in my own words

OpenStudy (anonymous):

How far does the tip of the minute hand travel during that time? the hand if 4 inches long multiply the result above by 4

OpenStudy (barrelracing):

ok im back.

OpenStudy (anonymous):

ok we are done right?

OpenStudy (barrelracing):

need help for the second please

OpenStudy (anonymous):

i wrote the answer above `

OpenStudy (barrelracing):

4 inches long?

OpenStudy (anonymous):

the angle is actually defined by the arc length over the radius your radius is 4 and your angle is \(\frac{2\pi}{3}\) therefore the arc length is the product of those two numbers

OpenStudy (barrelracing):

divide 2pi/3 by 4 to get the arc length?

OpenStudy (anonymous):

"product"

OpenStudy (anonymous):

aka multiply

OpenStudy (barrelracing):

8.37?

OpenStudy (anonymous):

i would just write \[4\times \frac{2\pi}{3}=\frac{8\pi}{3}\] and leave it at that

OpenStudy (barrelracing):

thank you so much i understand now

OpenStudy (anonymous):

yw